Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments—Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 13,750 $ 17,250 $ 31,000
Estimated variable manufacturing overhead per machine-hour $ 2.90 $ 3.70
Job P Job Q
Direct materials $ 28,000 $ 15,500
Direct labor cost $ 33,000 $ 13,500
Actual machine-hours used:
Molding 3,200 2,300
Fabrication 2,100 2,400
Total 5,300 4,700

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1–8, assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. For questions 9–15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

5. What was the total manufacturing cost assigned to Job Q? (Do not round intermediate calculations.)

5a. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

5b. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis assuming 20 units were produced for Job P and 30 units were produced for Job Q? (Do not round intermediate calculations. Round your final answers to nearest whole dollar.)

5c. What was Sweeten Company’s cost of goods sold for March? (Do not round intermediate calculations.)

5d. What were the company’s predetermined overhead rates in the Molding Department and the Fabrication Department? (Round your answers to 2 decimal places.)
Answer #1
Estimated variable manufacturing overhead 12800 =(2500*2.9)+(1500*3.7)
Estimated total fixed manufacturing overhead 31000
Estimated total manufacturing overhead 43800
Divide by Estimated total machine-hours used 4000
Plantwide predetermined overhead rate 10.95 per MH
Job P
Direct materials 28000
Direct labor cost 33000
Manufacturing overhead applied 58035 =5300*10.95
Total manufacturing cost Job P 119035
5
Job Q
Direct materials 15500
Direct labor cost 13500
Manufacturing overhead applied 51465 =4700*10.95
Total manufacturing cost Job Q 80465
5a
Total manufacturing cost Job Q 80465
Divide by units 30
Unit product cost Job Q 2682
5b
Job P Job Q
Total manufacturing cost 119035 80465
Add: Markup @ 80% 95228 64372
Total price for the job 214263 144837
Divide by units 20 30
Selling price per unit 10713 4828
5c
Cost of goods sold for March 199500 =119035+80465
5d
Predetermined overhead rates:
Molding Department 8.40 =2.9+(13750/2500)
Fabrication Department 15.20 =3.7+(17250/1500)